OpenStudy (anonymous):

I need help on Canceling out division. I know u have to undo it by multiplying it to the other side of the equation. Like n/7=11 n=77. But how would u do it on like for example n/7=x/11. Do you multiply 7 to both x and 11 or just 11??

OpenStudy (whpalmer4):

\[\frac{n}{7} = \frac{x}{11}\]Multiply both sides by 7: \[\frac{7n}{7} = \frac{7x}{11}\]Multiply both sides by 11: \[\frac{7*11n}{7} = \frac{11*7x}{11}\]Cancel common factors \[11n=7x\]

OpenStudy (whpalmer4):

If you multiply both sides of the equation by the product of all of the denominators, all the fractions go away. That may be a larger number than you need to use, but it will always work. The smallest number that will work is the LCM or least common multiple of all the denominators.

OpenStudy (anonymous):

So u only mutiplu it to the numerators not denominators ?

OpenStudy (whpalmer4):

In that particular equation, you could also solve by cross multiplying: \[\frac{n}{7} = \frac{x}{11}\]\[11n = 7x\] Multiplying the numerators is sufficient. Just remember to multiply everything!